OpenStudy (anonymous):

julie and eric are going a constant speed down stream 24 miles for 4 hours at the same rate they go upsteam 24 miles for 6 hours, what is the rate of the current?

OpenStudy (anonymous):

Which formula will we use? d = ?

OpenStudy (anonymous):

im thinking d=rt

OpenStudy (anonymous):

Yes, now r can get complicated .... upstream we are going x - y and down stream we are going x + y does this make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the first equation is 24 = 4(x + y) can you get the second equation? 24 = ???

OpenStudy (anonymous):

24=6(x-y)

OpenStudy (anonymous):

so 4(x +y)=6(x-y)?

OpenStudy (anonymous):

Yes, but I do not think we can solve this with x's and y's how about 24 = 4x + 3y 24 = 6x - 6y and multiply the first equation by 2 to obtain 48 = 8x + 6y 24 = 6x - 6y and then eliminate the y's to obtain 72 = ??? what is x???

OpenStudy (anonymous):

i think the current is 1 mph

OpenStudy (anonymous):

Do you see that 72 = 14x?